OpenStudy (xochital):

Consider the line that passes through the point (3, -6) and has a slope of 4. Part 1: Write the equation of this line using point-slope form. Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work.

OpenStudy (xochital):

what i have so far.... Part 1: y + 6 = 4(x - 3) Part 2: distribute the 4 y = 4x - 12 i don't know what to do for part 3.

OpenStudy (anonymous):

Standard form: \[Ax+By=C\]

OpenStudy (xochital):

right so would it be... 4x + 12y = C ?

OpenStudy (anonymous):

Not quite, no. Move all the terms with x and y to one side first.

OpenStudy (xochital):

4x - y = 12?

OpenStudy (anonymous):

Yep, that's it

OpenStudy (xochital):

thank you so much! :)

OpenStudy (tkhunny):

What? Where did the "+6" go? It was hanging out with the 'y', there, for a moment. Then, it just disappeared. @xochital

OpenStudy (anonymous):

Good eye, @tkhunny